# Singular flat bands in three dimensions: Landau level spreading, quantum geometry, and Weyl reconstruction

### v1

@@ A.1 $k_{z}\rightarrow\infty$ regime
$\vec{\Phi}_{+1}=\frac{1}{\sqrt{2\lambda_{j}(2j+1)}}\left(\begin{array}[]{c}\sqrt{j}k_{z}l_{B}\\ 2j+1\\ \sqrt{j+1}k_{z}l_{B}\end{array}\right)$ (68)
@@
where $[F]_{s,s^{\prime}}=\vec{\Phi}_{s}^{\dagger}H_{\mathrm{Z}}\vec{\Phi}_{s^{\prime}}$ with $s,s^{\prime}=\pm 1$ and $D=\hbar\omega_{B}\lambda+\vec{\Phi}_{0}^{\dagger}H_{\mathrm{Z}}\vec{\Phi}_{0}$ represent the blocks of flat and dispersive bands, and $U=(\vec{\Phi}_{0}^{\dagger}H_{\mathrm{Z}}\vec{\Phi}_{-1},\vec{\Phi}_{0}^{\dagger}H_{\mathrm{Z}}\vec{\Phi}_{+1})$ is coupling between two sectors. For $\lambda\gg 1$, the two blocks are energetically separated, and the Hamiltonian in the $s=\pm 1$ sector describes the low-energy physics. Applying standard second-order perturbation theory, the projected Hamiltonian in this sector is given by
@@
±$\pm\sqrt{\left(2+\frac{k_{z}^{2}l_{B}^{2}}{\lambda}-\frac{j(j+1)}{\lambda^{2}}\right)^{2}+\frac{32k^{2}_{z}l_{B}^{2}j(j+1)}{\lambda}}\Bigg{)}.$ (75)
Here, we use the estimations $k_{z}\l_{B}\lesssim\sqrt{\lambda}$ and $j\lesssim\lambda$, retaining terms up to $\lambda^{-1}$ while neglecting higher-order correction of $\mathcal{O}(\lambda^{-2})$ The expression Eq. (A.1) captures the asymptotic behavior in Fig. 3. In the more restrictive limit $j\ll k_{z}l_{B}$, the expression reduces to Eq. (39).
@@ Appendix B Zeeman splitting in Kagome Lattice model
so the eigenvalue $j>-1$ of $J^{(\mathrm{Kag})}_{z}$ is good quantum number. The projected Hamiltonian in $j$-sector for $j\geq 0$ is given by
@@
For $j=-1$ sector, the Hamiltonian is represented by the lower-right element of $H_{LL}^{(\mathrm{Kag},-1)}$, yielding a Landau level at $E=\hbar\omega_{B}/4$. The first term in Eq. (93) clearly gives eigenvalues
@@ Appendix C Formulating the density of states under the magnetic field
where $\Lambda$ represents the system size and the Landau level degeneracy is given by $g_{B}=\Lambda^{2}/l_{B}^{2}$. The occupation of the states is determined by the Fermi distribution function $f(E)=(1+e^{\beta E})^{-1}$, which is centered at the Fermi energy. We subtract $1/d^{3}$ in Eq. (96) to define the zero point of the charge density, $\rho=0$ to correspond to a fully filled flat band, in which two electrons occupy each unit cell of volume $2d^{3}$ (see Fig. 2(a)). To ensure that the total state density, accounting for three orbitals per unit cell, satisfies the condition $\lim_{E\rightarrow\infty}\rho(E,B)=1/(2d^{3})$, we introduce a cutoff for the Landau level index $n=j-m$ as $n_{c}=l_{B}^{2}/d^{2}$. Using the charge density from Eq. (96), the density of states at the Fermi energy is given by
@@
[…]
@@
$B_{\rho}=\frac{\hbar}{e}|\rho|^{2/3}.$ (100)
Using this scale, normalized Fermi energy in Eq. (99) is given by $\tilde{E}_{\mathrm{F}}=\tilde{E}_{\mathrm{F}}(B/B_{\rho})$ Therefore, the normalized density of states is expressed as
@@ Appendix D Divergent magnetic response in the flat-band limit
[…]
@@
which qualitatively reproduces the linear behavior in Fig. 11 (b). The solution Eq. (104) is plotted as the dashed line in Fig. 11 (b) which shows the order of the slope of $\tilde{u}$. However, analysis quantitatively overestimate the slope of $u-B$ curve. This is because around $j\sim n_{c}$ in the integration interval $|k_{z}l_{B}|\lesssim\sqrt{n_{c}}$ the region where asymptotic behavior in Eq. (39) is not relevant becmes non-negligible. Nontheless, the linear dependence of $\tilde{u}$ remains robust, as confirmed by the full numerical calculation.
@@
Figure 11: Numerically obatained field dependence of the total energy density $u$ (Eq. (LABEL:eq:utot)), blue and red dots in log-log plot (a) are the result for different fillings: $\nu=1$ where the half filling of twofold degenerate flat band, $\nu=2$ where the two flat bands are fully filled. (b) and (c) is the linear scale plot for the two cases.
@@ Appendix E Matrix Representation of Angular momentum
Appendix F Artificially enhanced Zeeman term
To clarify the correspondence between picture of the chiral Landau levels shown in Fig. 9 and real Landau level Fig. 8(a), we consider a modified effective Hamiltonian where the Zeeman term is artificaially enhanced,
$H_{\mathrm{eff+\eta Z}}(\bm{k})=H_{\mathrm{eff}}(\bm{k})+\eta H_{\mathrm{Z}}\$ (113)
and its Landau-level version,
$H_{\mathrm{B},\eta}(j,k_{z})=\tilde{H}_{B}(j,k_{z})+\eta H_{\mathrm{Z}},$ (114)
using Eqs. (49),  (52), and (56). While this enhancement amplifies the Zeeman splitting, it does not change the topological properties of the system. Instead, it reduces the Landau level spacing relative to the Weyl points separation, allowing for a clearler comparison between discrete spectrum and continuum Weyl physics.
To illustrate this behavior, we examine the case of $\ell=3/2$ with $\eta=100$, as shown in Fig. 12. While Zeeman splitting is substantially enhanced, the spectrum in Fig. 12 is related to that in Fig. 8(a) via the rescaling $E\rightarrow E/\eta$ and $k_{z}\rightarrow k_{z}/\sqrt{\eta}$. In this rescaled system, the Landau level spacing becomes relatively smaller compared to the band structure, making the underlying features more apparent. The two Weyl points along the $k_{z}$-axis in Fig. 12(a) arises from a band inversion between states with angular momentum projections: one between $m_{z}=-1/2$ and $m^{\prime}_{z}=-3/2$, and the other between $m_{z}=+1/2$ and $m^{\prime}_{z}=-3/2$. Owing to the correspondence between the Chern number and the total angular momentum projection $m_{z}$ along the high symmetry axis, the helicities of these Weyl points are given by $\chi=m_{z}-m_{z}^{\prime}$, resulting in $\chi=1$ and $2$, respectively.
In the same manner as for $\ell=1$ in Fig. 4, $\ell=3/2$ system exhibits a single flat chiral Landau level associated with the $\chi=1$ Weyl point. This modes corresponds to $j=-3/2$ state, shown as the red line in Fig. 12(b). Additionally, at the $\chi=2$ Weyl points located at larger $k_{z}$, the $j=-1/2$ mode (yellow curve in Fig. 12(b)) approaches and aligns with the $j=-3/2$ mode. Togather, they form two chiral Landau levels that connect states below and above the Weyl points.

### v2

@@ A.1 $k_{z}\rightarrow\infty$ regime
$\vec{\Phi}_{+1}=\frac{1}{\sqrt{2\lambda_{j}(2j+1)}}\left(\begin{array}[]{c}\sqrt{j}k_{z}l_{B}\\ 2j+1\\ -\sqrt{j+1}k_{z}l_{B}\end{array}\right)$ (72)
@@
where $F]_{s,s^{\prime}}=\vec{\Phi}_{s}^{\dagger}H_{\mathrm{Z}}\vec{\Phi}_{s^{\prime}}$ with $s,s^{\prime}=\pm 1$ and $D=\hbar\omega_{B}\lambda+\vec{\Phi}_{0}^{\dagger}H_{\mathrm{Z}}\vec{\Phi}_{0}$ represent the blocks of flat and dispersive bands, and $U=(\vec{\Phi}_{0}^{\dagger}H_{\mathrm{Z}}\vec{\Phi}_{-1},\vec{\Phi}_{0}^{\dagger}H_{\mathrm{Z}}\vec{\Phi}_{+1})$ is coupling between two sectors. For $\lambda\gg 1$, the two blocks are energetically separated, and the Hamiltonian in the $s=\pm 1$ sector describes the low-energy physics. Applying standard second-order perturbation theory, the projected Hamiltonian in this sector is given by
@@
$\pm\sqrt{\left(2+\frac{k_{z}^{2}l_{B}^{2}}{\lambda}-\frac{j(j+1)}{\lambda^{2}}\right)^{2}+\frac{32k^{2}_{z}l_{B}^{2}j(j+1)}{\lambda}}\Bigg).$ (79)
Here, we use the estimations $k_{z}\l _{B}\lesssim\sqrt{\lambda}$ and $j\lesssim\lambda$, retaining terms up to $\lambda^{-1}$ while neglecting higher-order correction of $\mathcal{O}(\lambda^{-2})$ . The expression Eq. (A.1) captures the asymptotic behavior in Fig. 3. In the more restrictive limit $j\ll k_{z}l_{B}$, the expression reduces to Eq. (39).
@@ Appendix B Zeeman splitting in Kagome Lattice model
so the eigenvalue j>1j>-1 of $J^{(\mathrm{Kag})}_{z}$ is good quantum number. The projected Hamiltonian in $j$-sector for $j\geq 0$ is given by
@@
For j=1j=-1 sector, the Hamiltonian is represented by the lower-right element of $H_{LL}^{(\mathrm{Kag},-1)}$, yielding a Landau level at $E=\hbar\omega_{B}/4$. The first term in Eq. (97) clearly gives eigenvalues
@@ Appendix C Formulating the density of states under the magnetic field
where $\Lambda$ represents the system size and the Landau level degeneracy is given by $g_{B}=\Lambda^{2}/l_{B}^{2}$. The occupation of the states is determined by the Fermi distribution function $f(E)=(1+e^{\beta E})^{-1}$, which is centered at the Fermi energy. We subtract $1/d^{3}$ in Eq. (100) to define the zero point of the charge density, $\rho=0$ to correspond to a fully filled flat band, in which two electrons occupy each unit cell of volume $2d^{3}$ (see Fig. 2(a)). To ensure that the total state density, accounting for three orbitals per unit cell, satisfies the condition $\lim_{E\rightarrow\infty}\rho(E,B)=1/(2d^{3})$, we introduce a cutoff for the Landau level index n=jmn=j-m as $n_{c}=l_{B}^{2}/d^{2}$. Using the charge density from Eq. (100), the density of states at the Fermi energy is given by
@@
$\tilde{\rho}(\tilde{E},\gamma)=\frac{1}{2\pi}\sum_{\begin{subarray}{c}\nu,j\\ n<\gamma^{-2}\end{subarray}}\int_{-\frac{\pi}{2\gamma}}^{\frac{\pi}{2\gamma}}d\tilde{k}\tilde{f}(\tilde{E}_{\nu}^{(j)}-\tilde{E})-\gamma^{-3}.$ (102)
@@
$B_{\rho}=\frac{\hbar}{e}|\rho|^{2/3},$ (104)
normalized Fermi energy in Eq. (103) is given by $\tilde{E}_{\mathrm{F}}=\tilde{E}_{\mathrm{F}}(B/B_{\rho})$ Therefore, the normalized density of states is expressed as
@@ Appendix D Divergent magnetic response in the flat-band limit
$\tilde{u}=\frac{\gamma^{5}}{2\pi}\sum_{\begin{subarray}{c}\nu,j\\ n<1/\gamma^{2}\end{subarray}}\int_{-\frac{\pi}{2\gamma}}^{\frac{\pi}{2\gamma}}d\tilde{k}\tilde{E}_{\nu}^{(j)}(\tilde{k})\tilde{f}(\tilde{E}_{\nu}^{(j)}(\tilde{k})-\tilde{E}_{F})$ (106)
@@
which qualitatively reproduces the linear behavior in Fig. 12 (b). The solution Eq. (108) is plotted as the dashed line in Fig. 12 (b) which shows the order of the slope of $\tilde{u}$. However, analysis quantitatively overestimate the slope of uBu-B curve. This is because around $j\sim n_{c}$ in the integration interval $|k_{z}l_{B}|\lesssim\sqrt{n_{c}}$ the region where asymptotic behavior in Eq. (39) is not relevant becmes non-negligible. Nontheless, the linear dependence of $\tilde{u}$ remains robust, as confirmed by the full numerical calculation.
@@
Figure 12: Numerically obatained field dependence of the total energy density $u$ (Eq. (106)), blue and red dots in log-log plot (a) are the result for different fillings: $\nu=1$ where the half filling of twofold degenerate flat band, $\nu=2$ where the two flat bands are fully filled. (b) and (c) is the linear scale plot for the two cases.
@@ Appendix E Matrix Representation of Angular momentum
Appendix F High-angular momentum model with enhanced Zeeman term
To clarify the correspondence between picture of the chiral Landau levels shown in Fig. 10 and real Landau level Fig. 9(a), we consider a modified effective Hamiltonian in which the Zeeman term is artificaially enhanced. This model is defined by Eq. (41) and (42) along with Eq. (52), (54), and (55). While the enhanced Zeeman term increases the energy splitting, it does not change the topological properties of the system. Instead, it reduces the Landau level spacing relative to the separation of the Weyl points. This allows for a clearler comparison between discrete spectrum and continuum Weyl physics, as also discussed in Sec. III.2.
To illustrate this behavior in higher angular momentum singular flat band model, we examine the case of $\ell=3/2$ with $\eta=100$, as shown in Fig. 13. While Zeeman splitting is substantially enhanced, the spectrum in Fig. 13 is related to that in Fig. 9(a) via the rescaling $E\rightarrow E/\eta$ and $k_{z}\rightarrow k_{z}/\sqrt{\eta}$. In this rescaled system, the Landau level spacing becomes relatively smaller compared to the band structure, making the underlying features more apparent. The two Weyl points along the $k_{z}$-axis in Fig. 13(a) arises from a band inversion between states with angular momentum projections: one between $m_{z}=-1/2$ and $m^{\prime}_{z}=-3/2$, and the other between $m_{z}=+1/2$ and $m^{\prime}_{z}=-3/2$. Owing to the correspondence between the Chern number and the total angular momentum projection $m_{z}$ along the high symmetry axis, the helicities of these Weyl points are given by $\chi=m_{z}-m_{z}^{\prime}$, resulting in $\chi=1$ and $2$, respectively.
In the same manner as for $\ell=1$ in Fig. 4, $\ell=3/2$ system exhibits a single flat chiral Landau level associated with the $\chi=1$ Weyl point. This modes corresponds to j=3/2j=-3/2 state, shown as the red line in Fig. 13(b). Additionally, at the $\chi=2$ Weyl points located at larger $k_{z}$, the j=1/2j=-1/2 mode (yellow curve in Fig. 13(b)) approaches and aligns with the j=3/2j=-3/2 mode. Togather, they form two chiral Landau levels that connect states below and above the Weyl points.